OpenStudy (anonymous):

calculate the modulus argument of -1+√3i

OpenStudy (amistre64):

the angle associated with arctan(-sqrt(3)/1)

OpenStudy (anonymous):

It is the solution of \[\sin \theta =\sqrt{3}/2\] and \[\cos \theta = -1/2\]